OpenStudy (faolan):

Fan N Medal pls help!! Which polynomial equation of least degree has -2, -2, 3, and 3 as four of its roots?

OpenStudy (faolan):

(x + 2)(x - 3) = 0 (x - 2)^-2(x + 3)^3 = 0 (x^-2 + 2)(x^2- 3) = 0 (x + 2)^2(x - 3)^2 = 0

OpenStudy (faolan):

@benlindquist @AloneS @Vestran01 @DannyBoyDan @-AngleWarrior-

OpenStudy (faolan):

@party_girl

OpenStudy (faolan):

thx for help......i totally understand it now